### v1

@@ 3 Results and Discussion
The CHROMIXS chromatograms indicated the presence of one single main peak eluting in a similar position for all the duplexes, with some small shoulders on either side of the peak (Supplementary Figure S1). All eluted peaks had 260:280nm UV absorbance ratios consistent with that of pure dsDNA (Supplementary Figure S1). Each dsDNA peak had some instability in predicted $R_{g}$ values across the eluted peak, possibly because of neighbouring peaks. However, frames were selected for analysis from regions where the $R_{g}$ was stable across consecutive frames within approximately 55\,Å or less (Supplementary Figure S1). In the case of GAGE6, the $R_{g}$ across the chosen frames varied from $45.3$ to 50.7850.78\,Å with an average of 48.0848.08\,Å. For GAGE6_1 the $R_{g}$ varied from 49.949.9\,Å to 52.952.9\,Å  with an average of 52.052.0\,Å. For GAGE6_2 the $R_{g}$ of the chosen frames ranged from 53.954.053.9-54.0\,Å with the $R_{g}$ in the surrounding frames varying from 51.554.051.5-54.0\,Å. In the case of GAGE6_3, the variation in the predicted $R_{g}$ across the chosen frames was minimal, ranging from 49.350.649.3-50.6\,Å  with an average of 50.150.1\,Å.
@@ 3.2 Sequence GAGE6
Figure 6: Sequence GAGE6: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
@@
Figure 6B shows the statistics of the bending angles $|\theta_{n}|$ and dihedral angles $\phi_{n}$ with their averages over conformations and standard deviations. For the bending angles this figure is analogous to Fig. 4 except that the conformations have been oriented as described in the Materials and Methods section. Only one of the double maxima that were forming a mirror image with respect to the middle of the sequence has subsisted, which confirms that the symmetric pattern was created through symmetry inherent in the SAXS measurement. The value $\Delta s/s$ marked on the figure corresponds to $(s^{\prime}-s)/s$, where $s$ and $s^{\prime}$ were the local sums of the bending angles in the two mirror domains used to orient the conformations. Their differences were significant, suggesting that one of the two mirror images was fictitious, which is confirmed by the figure as one of them has been fully erased in the orientation step.
@@
The three-dimensional plot of Fig. 6C shows the probability distribution of the bending angles $|\theta|$ for each site $n$ along the sequence. These distributions are normalised
$\int_{0}^{\pi}H(n,\theta)=1\quad{\mathrm{for~all~}}n$ (5)
@@
which allows us to calculate the fraction $W$ of conformations which have a bending angle above some value in a given domain. The result $W=0.06557$ is shown in Fig 6C for $35\leq n\leq 49$ and a lower value of $\theta$ of $45^{\circ}$. Thus, although most of the bending angles are well below $30^{\circ}$, in some regions of the sequence there are about $6\%$ of the conformations which show bending above $45^{\circ}$. The statistics clearly show that one region of the sequence is special, with a much higher flexibility and larger bending than elsewhere. The colour view of the three-dimensional histogram Fig. 6C, (see Supplementary material Fig. S8) suggests that the large bending angles for $35\leq n\leq 49$ are separated from the bulk of smaller bending angles, as if they result from some permanent bend in this region and not simply from large amplitude thermal fluctuations. As this part of the histogram only concerns $6\%$ of the conformations, the resolution of the analysis of the SAXS data does not however allow us to make a definitive statement on this point.
@@
We would like to stress that these statistical results which provide precise information on the local properties of the sample, are extracted from a probability distribution function which looks rather featureless although it results from a large ensemble of pair distances. It is remarkable that this smooth looking, almost straight function, actually encodes so much information. Now that we have oriented the conformations we can also compare the statistical results with the actual sequence of the sample. There is still an unknown point: should the sequence listed in Fig. 1 be oriented from 1 to 60 or from 60 to 1 ?
@@
Figure 7: Sequence GAGE6_1: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_1: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_1 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_1: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
@@ 3.3 Sequences GAGE6_1, GAGE6_2, GAGE6_3
Figure 7C confirms the presence of large bending angles in a narrow domain and the integration of $H(n,\theta)$ for $41\leq n\leq 52$ and $\theta>45^{\circ}$ shows that the fraction $W$ of bending angles above $45^{\circ}$ has drastically increased from 6.4% to 12.6%. The GAGE6_1 sequence therefore has a high probability of being sharply bent in a narrow region. It is interesting to notice the qualitative analogy between the experimental pair-distance distribution function for GAGE6_1 and the double-bump simulated $P(r)$ (Fig. 2) obtained with an atomistic model of sharply bent DNA. This strengthens the notion that $P(r)$ has a high sensitivity to molecular conformations, particularly bending, with sharp bending generating a function with two bumps.
Figure 8: Sequence GAGE6_2: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_2: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_2 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_2: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
@@
Figure 8B shows that the distribution of large bending angles along the sequence, as well as the domains where the dihedral angles have large standard deviations, have qualitatively changed. Figure 8B appears to be very consistent with the sequence of GAGE6_2. Its largest AT domains are $10\leq n\leq 13$ and a large domain from $n=50$ to $n=57$ which is only interrupted by a single GC pair at position $n=55$. This agrees well with the small increase of bending and twist fluctuations seen around $n=10$ in Fig. 8B and the larger bending angles and twist fluctuations in the range $44\leq n\leq 57$ split into two peaks by a dip in sites $52,53$ which could correspond to the GC pair at site $55$ within the resolution given by the analysis of the SAXS data. The bending angles $|\theta_{n}(i)|$ in this domain are however much smaller than in the sharp bending domain of GAGE6_1. This qualitative change from GAGE6_1 may nevertheless seem surprising if we think that the only difference in sequence between GAGE6_1 and GAGE6_2 is the elimination of an AT-rich domain between positions $38\leq n\leq 44$. Both GAGE6_1 and GAGE6_2 have a large AT-rich domain ($50\leq n\leq 57$) but in GAGE6_2 this region causes a smaller bending effect than GAGE6_1. This points out again that, although sequence and flexibility are correlated, the preferred conformations of a DNA sequence in solution are not only determined by the local sequence but collective effects at larger distances are also important. The histograms of the bending angles shown in Fig. 8C and Supplementary Fig. S8 are somewhat similar to the data from Fig. 7: the fraction of large bending angles in the range $44\leq n\leq 57$ is of the order of 9.6%, suggesting that GAGE6_2 is also sharply bent in a narrow region.
Figure 9: Top panel: sequence GAGE6_2. Comparison of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ (small circle points and thin black curve with error bars) and $P(r)$ given by the 1000 best conformations of the polymer model (thick red curve). The gray region for r<25r<25\;Å is the low-r domain in which the model cannot describe the internal structure of the double helix. Bottom panel: Statistics of the local bending angles $\theta_{n}(i)$ over the 1000 best model conformations: the red full line shows the average over $i$ of $|\theta_{n}(i)|$. The thin black line without symbols shows the same average limited to the bending angles $\theta$ above $5^{\circ}$ and dashed blue curve shows the standard deviations of $\theta_{n}(i)$. The thin black line without symbols shows the average of the dihedral angles and the thin black-dot-line shows their standard deviation. The vertical dashed lines show the boundaries of the domain where large bending was observed in the analysis of AutoGNOM-5-$P_{\mathrm{exp}}(r)$.
@@
In GAGE6_3, the last of the series of sequences derived from GAGE6, the last large AT-rich domain that spanned positions $50\leq n\leq 54$ has been replaced entirely by GC base pairs. Figure 10A shows that $P_{\mathrm{exp}}(r)$ has lost its two-lump structure and conforms in comparison to the other experimental functions more to the simulated shape of a straight duplex (Figure 2). The disagreement between the polymer model fit and $P_{\mathrm{exp}}(r)$ at large values of $r$ here may be a result of the former not accounting for hydration shell effects whilst the latter does. Although the large AT-rich regions of the $n>30$ range that we used to orient the conformations for previous sequences are no longer present, the sequence still has asymmetry. The domain $10\leq n\leq 26$ has several short AT sequences while its mirror image $35\leq n\leq 49$ now only has three AT pairs and is likely to be less flexible, being systematically enriched for GC content. This is confirmed by the sum $s$ and $s^{\prime}$ of $|\theta_{n}|$ in these two regions which differ systematically from each other and can be used to orient the conformations with respect to the sequence. Figure 10B, which does not display the symmetry expected for unoriented SAXS data, shows that the orientation procedure described in the Material and Methods section is again valid for GAGE6_3. Figure 10B shows a broad domain with large bending angles with a sharp maximum at position $n=26$ which coincides with two AT pairs in the middle of a GC sequence. The three-dimensional plot of the histograms $H(n,\theta)$ plotted in Fig. 10C and Supplementary Fig. S8 shows larger bending angles in the domain $17\leq n\leq 35$ but the integrated weight of the histograms for angles exceeding $45^{\circ}$ in this region shows that only 3.8% of the conformations exhibit large bending. As suggested by its sequence, mostly GC base pairs and only a few small domains with AT pairs, not wider than three consecutive sites, the GAGE6_3 sequence has mostly straight conformations.
@@
Figure 10: Sequence GAGE6_3: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_3: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_3 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_3: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.

### v2

@@ 3 Results and Discussion
3.1 Small Angle X-ray Scattering data
The chromatograms computed with the program CHROMIXS [CHROMIXS] indicated the presence of one single main peak eluting in a similar position for all the duplexes, with some small shoulders on either side of the peak (Supplementary figure S1). All eluted peaks had 260:280nm UV absorbance ratios consistent with that of pure dsDNA (Supplementary figure S1). Each dsDNA peak had some instability in predicted $R_{g}$ values across the eluted peak, possibly because of neighbouring peaks. However, frames were selected for analysis from regions where the $R_{g}$ was stable across consecutive frames within approximately 55\,Å or less (Supplementary figure S1). In the case of GAGE6, the $R_{g}$ across the chosen frames varied from $45.3$ to 50.7850.78\,Å with an average of 48.0848.08\,Å. For GAGE6_1 the $R_{g}$ varied from 49.949.9\,Å to 52.952.9\,Å with an average of 52.052.0\,Å. For GAGE6_2 the $R_{g}$ of the chosen frames ranged from 53.954.053.9-54.0\,Å with the $R_{g}$ in the surrounding frames varying from 51.554.051.5-54.0\,Å. In the case of GAGE6_3, the variation in the predicted $R_{g}$ across the chosen frames was minimal, ranging from 49.350.649.3-50.6\,Å with an average of 50.150.1\,Å.
The calculation of the molecular weight of the frames across the chromatography peak worked well for GAGE6_1, GAGE6_2, and GAGE6_3; however, it was unsuccessful for GAGE6, perhaps due to the fact that this experiment is slightly noisier than the others. As shown in Supplementary figure S2, the molecular weight of the frames across the peaks are in agreement with the expected molecular weight of a dsDNA duplex of this size ($\approx 37$kDa). This demonstrates that at least in the case of 3 of the 4 chromatograms, the samples are clearly monodisperse. Despite this calculation not working for the GAGE6 sample, the $R_{g}$ values of the chosen frames for the GAGE6 experiment are in close agreement, demonstrating that this sample is also likely to be monodisperse.
@@ 3.2 Sequence GAGE6
Figure 6: Sequence GAGE6: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$. The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation. (C) Sequence GAGE6: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
@@
Figure 6B shows the statistics of the bending angles $|\theta_{n}|$ and dihedral angles $|\varphi_{n}|$ with their averages over conformations and theit standard deviations. For the bending angles this figure is analogous to Fig. 4 except that the conformations have been oriented as described in the Materials and Methods section. Only one of the double maxima that were forming a mirror image with respect to the middle of the sequence has subsisted, which confirms that the symmetric pattern was created through symmetry inherent in the SAXS measurement. The value $\Delta s/s$ marked on the figure corresponds to $(s^{\prime}-s)/s$, where $s$ and $s^{\prime}$ were the local sums of the bending angles in the two mirror domains used to orient the conformations. Their differences were significant, suggesting that one of the two mirror images was fictitious, which is confirmed by the figure as one of them has been fully erased in the orientation step.
@@
The three-dimensional plot of Fig. 6C shows the probability $P(n,|\theta|)$ of the bending angles $|\theta_{n}|$ for each site $n$ along the sequence. These distributions are normalised
$\int_{0}^{\pi}P(n,|\theta|)\;\mathrm{d}|\theta|=1\quad{\mathrm{for~each~}}n$ (5)
@@
which allows us to calculate the fraction $W$ of the bending angles $|\theta_{n}|$ above a some value in a given domain. The result $W=0.06557$ is shown in Fig 6C for $35\leq n\leq 49$ and a lower value of $\theta$ of $45^{\circ}$. Thus, although most of the bending angles are well below $30^{\circ}$, in some regions of the sequence there are about $6\%$ of the bending angles above $45^{\circ}$. The statistics clearly show that one region of the sequence is special, with a much higher flexibility and larger bending than elsewhere. The colour view of the three-dimensional probability $P(n,|\theta|)$ (Fig. 6C), (see Supplementary material Fig. S9) suggests that the large bending angles for $35\leq n\leq 49$ are separated from the bulk of smaller bending angles, as if they result from some permanent bend in this region and not simply from large amplitude thermal fluctuations. As this part of the probability distribution only concerns $6\%$ of the bending angles, the resolution of the analysis of the SAXS data does not however allow us to make a definitive statement on this point.
@@
We would like to stress that these statistical results which provide precise information on the local properties of the sample, are extracted from a probability distribution function $P_{\mathrm{exp}}(r)$ which looks rather featureless although it results from a large ensemble of pair distances. It is remarkable that this smooth looking, almost straight function, actually encodes so much information. Now that we have oriented the conformations we can also compare the statistical results with the actual sequence of the sample. There is still an unknown point: should the sequence listed in Fig. 1 be oriented from 1 to 60 or from 60 to 1 ?
@@
Figure 7: Sequence GAGE6_1: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_1: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$. The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation. (C) Sequence GAGE6_1: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
@@ 3.3 Sequences GAGE6_1, GAGE6_2, GAGE6_3
Figure 7C confirms the presence of large bending angles in a narrow domain and the integration of $P(n,|\theta|)$ for $41\leq n\leq 52$ and $\theta>45^{\circ}$ shows that the fraction $W$ of bending angles $|\theta_{n}|$ above $45^{\circ}$ has drastically increased from 6.4% to 12.6%. The GAGE6_1 sequence therefore has a high probability of being sharply bent in a narrow region. It is interesting to notice the qualitative analogy between the experimental pair-distance distribution function for GAGE6_1 and the double-bump simulated $P(r)$ (Fig. 2) obtained with an atomistic model of sharply bent DNA. This strengthens the notion that $P(r)$ has a high sensitivity to molecular conformations, particularly bending, with sharp bending generating a function with two bumps.
Figure 8: Sequence GAGE6_2: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_2: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$. The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation. (C) Sequence GAGE6_2: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
@@
Figure 8B shows that the distribution of large bending angles along the sequence, as well as the domains where the dihedral angles have large standard deviations, have qualitatively changed. Figure 8B appears to be very consistent with the sequence of GAGE6_2. Its largest AT domains are $10\leq n\leq 13$ and a large domain from $n=50$ to $n=57$ which is only interrupted by a single GC pair at position $n=55$. This agrees well with the small increase of bending and twist fluctuations seen around $n=10$ in Fig. 8B and the larger bending angles and twist fluctuations in the range $44\leq n\leq 57$ split into two peaks by a dip in sites $52,53$ which could correspond to the GC pair at site $55$ within the resolution given by the analysis of the SAXS data. The bending angles $|\theta_{n}(i)|$ in this domain are however much smaller than in the sharp bending domain of GAGE6_1. This qualitative change from GAGE6_1 may nevertheless seem surprising if we think that the only difference in sequence between GAGE6_1 and GAGE6_2 is the elimination of an AT-rich domain between positions $38\leq n\leq 44$. Both GAGE6_1 and GAGE6_2 have a large AT-rich domain ($50\leq n\leq 57$) but in GAGE6_2 this region causes a smaller bending effect than GAGE6_1. This points out again that, although sequence and flexibility are correlated, the preferred conformations of a DNA sequence in solution are not only determined by the local sequence but collective effects at larger distances are also important. The probability distribution of the bending angles $P(n,|\theta|)$ shown in Fig. 8C and Supplementary Fig. S9 are somewhat similar to the data from Fig. 7: the fraction of large bending angles $|\theta_{n}|$ in the range $44\leq n\leq 57$ is of the order of 9.6%, suggesting that GAGE6_2 is also sharply bent in a narrow region.
@@
Figure 9: Top panel: sequence GAGE6_2. Comparison of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ (small circle points and thin black curve with error bars) and $P(r)$ given by the 1000 best conformations of the polymer model (thick red curve). The gray region for r<25r<25\;Å is the low-r domain in which the model cannot describe the internal structure of the double helix. Bottom panel: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$. The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation. The vertical dashed lines show the boundaries of the domain where large bending was observed in the analysis of AutoGNOM-5-$P_{\mathrm{exp}}(r)$.
@@
In GAGE6_3, the last of the series of sequences derived from GAGE6, the last large AT-rich domain that spanned positions $50\leq n\leq 54$ has been replaced entirely by GC base pairs. Figure 10A shows that $P_{\mathrm{exp}}(r)$ has lost its two-lump structure and conforms in comparison to the other experimental functions more to the simulated shape of a straight duplex (Figure 2). The disagreement between the polymer model fit and $P_{\mathrm{exp}}(r)$ at large values of $r$ here may be a result of the former not accounting for hydration shell effects whilst the latter does. Although the large AT-rich regions of the $n>30$ range that we used to orient the conformations for previous sequences are no longer present, the sequence still has asymmetry. The domain $10\leq n\leq 26$ has several short AT sequences while its mirror image $35\leq n\leq 49$ now only has three AT pairs and is likely to be less flexible, being systematically enriched for GC content. This is confirmed by the sum $s$ and $s^{\prime}$ of $|\theta_{n}|$ in these two regions which differ systematically from each other and can be used to orient the conformations with respect to the sequence. Figure 10B, which does not display the symmetry expected for unoriented SAXS data, shows that the orientation procedure described in the Material and Methods section is again valid for GAGE6_3. Figure 10B shows a broad domain with large bending angles with a sharp maximum at position $n=26$ which coincides with two AT pairs in the middle of a GC sequence. The three-dimensional plot of the probability $P(n,|\theta|)$ plotted in Fig. 10C and Supplementary Fig. S9 shows larger bending angles in the domain $17\leq n\leq 35$ but the integrated weight of the probability $P(n,|\theta|)$ for angles exceeding $45^{\circ}$ in this region shows that only 3.8% of the angles correspond to large bending. As suggested by its sequence, mostly GC base pairs and only a few small domains with AT pairs, not wider than three consecutive sites, the GAGE6_3 sequence has mostly straight conformations.
@@
Figure 10: Sequence GAGE6_3: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_3: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$. The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation. (C) Sequence GAGE6_3: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.